# Variations on five-dimensional sphere packings

### v2

@@ 1. Introduction
Figure 1.1. The centers of the spheres in the face centered cubic kissing arrangement on the left, and the centers of the spheres in the hexagonal close packing kissing arrangement on the right. To produce the arrangement on the right from the one on the left, the top layer is deleted and the bottom layer is reflected across the central layer (the hexagon).
@@ 2. Five-dimensional kissing configurations
The kissing number in five dimensions appears to be $40$, although the best upper bound that has been proved is $44$ (from [15]). The first construction achieving $40$ is implicit in Korkine and Zolotareff’s 1873 paper [9], where they constructed the $D_{5}$ root lattice. Its root system achieves a kissing number of $40$ as the permutations of the points (±$(\pm 1,\pm 1,0,0,0)$; these points form a kissing configuration because they each have squared norm $2$ and the inner product between distinct points is always at most $1$.
In 1967, Leech [11] constructed a different kissing configuration of the same size, not isometric to the $D_{5}$ root system. He split $D_{5}$ into four-dimensional layers based on the fifth coordinate: there are $8$ points with fifth coordinate $1$, $24$ with 00, and $8$ with $-1$. The top layer consists of the points
(±$(\overbrace{\pm 1,0,0,0}^{\text{permute}},1),$
@@
(±$(\overbrace{\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2}}^{\text{odd \# of $-$ signs}},1)$
yields a different kissing configuration $L_{5}$, which breaks the reflection symmetry across the central hyperplane with fifth coordinate 00. One could also use an even number of minus signs, or replace both the top and bottom layers, but these modifications are all isometric to $D_{5}$ or $L_{5}$. For comparison, such a construction cannot be carried out in four dimensions, since the optimal kissing configuration is unique in four dimensions [6].
@@
(±$(\overbrace{\pm 1,\pm 1,0,0}^{\text{permute}},0)$
@@
(±$(\overbrace{\pm 1,0,0,0}^{\text{permute}},0),\quad(\overbrace{\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2}}^{\text{odd \# of $-$ signs}},0),\quad\text{and}\quad(\overbrace{\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2}}^{\text{even \# of $-$ signs}},0).$
@@
cyclically permutes these cross polytopes in four dimensions, so they all play the same role. Up to isometry, there are only two kissing configurations that can be obtained in this way: $D_{5}$, which has reflection symmetry across the hyperplane, and $L_{5}$, which does not. One can see that they are distinct by counting antipodal points: $D_{5}$ is closed under multiplication by $-1$, while $L_{5}$ is not. Instead, only $24$ of the points in $L_{5}$ come in antipodal pairs, namely those in $D_{4}$.
In 2023, Szöllősi [20] found a new kissing configuration $Q_{5}$ via a computer search. It can be described similarly to $L_{5}$, but using different cross sections: the $D_{5}$ root system contains $20$ vectors with a coordinate sum of 00 and $10$ each with coordinate sums of ±$\pm 2$. Here the central cross section is the $A_{4}$ root system. This cross section is suboptimal as a four-dimensional kissing configuration, but it nevertheless extends to a seemingly optimal configuration in five dimensions. The layer with coordinate sum $2$ consists of the antipodes of the points with coordinate sum $-2$, but it is not the reflection of that layer across the hyperplane with coordinate sum 00. Reflection across this hyperplane consists of the map
@@
[…]
Our new configuration, which we call $R_{5}$, can be obtained as follows by modifying $L_{5}$. The $L_{5}$ configuration behaves similarly to $D_{5}$ with respect to coordinate sums: there are $20$ points with coordinate sum 00 and $10$ each with ±$\pm 2$. The central cross section is no longer the $A_{4}$ root system, but rather a modification that replaces four adjacent points with a different regular tetrahedron. Again we can replace one of the adjacent layers with the other’s reflection, to obtain $R_{5}$ from $L_{5}$. Only $12$ of the points in $R_{5}$ come in antipodal pairs, and so it cannot be isometric to any of the three other configurations. This yields the following theorem:
@@
Configuration $-1$ $-4/5$ $-3/4$ $-1/2$ $-3/10$ $-1/4$ 00 $1/5$ $1/2$
$D_{5}$ $20$ 00 00 $240$ 00 00 $280$ 00 $240$
$L_{5}$ $12$ 00 $32$ $192$ 00 $32$ $272$ 00 $240$
$Q_{5}$ $10$ $30$ 00 $180$ $60$ 00 $250$ $10$ $240$
$R_{5}$ $6$ $30$ $20$ $144$ $60$ $28$ $242$ $10$ $240$
Table 2.2. The four five-dimensional kissing configurations with $40$ points. The subdivisions in the list are $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$, respectively, and the points with fractional coordinates are the only ones that differ from those in $D_{5}$.
@@
Table 2.1 lists the inner products that occur in these configurations and their multiplicities, while Table 2.2 lists coordinates for the points themselves. The symmetry groups of $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$ have orders $3840$, $384$, $240$, and $48$, respectively. Specifically, the symmetries of $D_{5}$ are the signed permutations of the coordinates, and those of $L_{5}$ are the signed permutations that fix the fifth coordinate. The symmetries of $Q_{5}$ are generated by the permutations of the coordinates and reflection across the hyperplane, and the symmetries of $R_{5}$ are generated by the permutations of the coordinates that fix the fifth coordinate and reflection across the hyperplane.
@@ 3. Six-dimensional kissing configurations
(±$(\pm\sqrt{2/5},\dots,\pm\sqrt{2/5})$
@@
If we start with $Q_{5}$ or $R_{5}$, it initially seems that the same approach might produce an extension to six dimensions, but it cannot be fully carried out. These configurations again have $32$ deep holes: those on one side of the hyperplane with coordinate sum 00 are the same as in $D_{5}$ or $L_{5}$, respectively, while those on the other side are the reflections across this hyperplane. However, continuing the construction requires a set of $16$ holes with maximal inner product $2/5$ between them, and one can check by depth-first search that no such set exists. We have not ruled out the possibility that some other construction might give a six-dimensional kissing configuration with $Q_{5}$ or $R_{5}$ as a cross section, but it seems unlikely:
@@ 4.1. Packings that fiber over $D_{3}$
To form a five-dimensional packing, we will use a four-colored point configuration in $\mathbb{R}^{2}$, by which we mean a discrete point set in $\mathbb{R}^{2}$ in which each point is labeled with the color 00, $1$, $2$, or $3$. We call two colors adjacent if they differ by $1$ modulo $4$, and opposite if they differ by $2$. If the point $v$ is colored $i$, then we center spheres of radius 2/$\sqrt{2}/2$ at the points in $\{v\}\times(D_{3}+t_{i})$. In order for the spheres not to overlap, the four-colored point configuration must satisfy the following constraints:
@@
A valid coloring is one that satisfies these constraints. One can check that if a valid four-colored point configuration has $\delta$ points per unit area in $\mathbb{R}^{2}$, then the corresponding sphere packing in $\mathbb{R}^{5}$ has packing density
$\frac{\pi^{2}\sqrt{2}}{30}\,\delta.$
In particular, matching the density of the $D_{5}$ root lattice amounts to achieving $\delta=1$. Presumably it is impossible to achieve $\delta>1$:
@@
To achieve $\delta=1$, we can tile the plane with two types of triangles: isosceles triangles with side lengths $\sqrt{5}/2,\sqrt{5}/2,1$ or $1,1,\sqrt{2}$, which we call $\Delta_{1}$ and $\Delta_{2}$. Both of these triangles have area $1/2$, and it follows that in every edge-to-edge tiling, there is one vertex per unit area. [^2] The angles in a triangle add up to $\pi$, which means they account for half of the total angle surrounding a single vertex. Therefore there must be half a vertex per triangle on average in any edge-to-edge tiling. Our five-dimensional sphere packings are obtained from valid four-colorings of the vertices of these tilings. Specifically, vertices at distance $\sqrt{2}$ have the same color, those are distance $\sqrt{5}/2$ have adjacent colors, and those at distance $1$ have opposite colors.
length $1$ (opposite colors)length $\sqrt{5}/2$ (adjacent colors)length $\sqrt{2}$ (same color)$\Delta_{1}$$\Delta_{2}$$q$$p$$a$$q$$p$$b$$c$$d$
@@
As mentioned in the introduction, the existence of more general packings such as Figure 4.2 disproves conjectures of Conway and Sloane [5] and Kuperberg [10]. We propose the following conjecture as a salvage for Conjecture 15 in [10].
@@ 4.2. Uniform packings
The uniform packing from Figure 4.2 with kissing configuration $Q_{5}$ is $2$-periodic, and the one with kissing configuration $R_{5}$ is $4$-periodic, with lattice bases and translation vectors shown in Table 4.2. Andreanov and Kallus [1] showed that no $2$-periodic sphere packing in $\mathbb{R}^{5}$ can be denser than the $D_{5}$ root lattice, and they obtained a complete list of the densest such packings. In the process they discovered the $Q_{5}$ uniform sphere packing, which also occurs in the acknowledgements of [20]. No such optimality theorem or classification is known beyond $2$-periodic packings. As for the remaining uniform packings, the $D_{5}$ root lattice is of course $1$-periodic, and the three $L_{5}$ uniform packings are $2$-periodic, $3$-periodic, and $4$-periodic (see [5]).
@@ 4.3. Symmetries of uniform packings
For the $Q_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{4}$ root lattice and a one-dimensional lattice (see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{5}\times C_{2}^{2}$, where the symmetric group $S_{5}$ permutes the five coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by $-1$ and the reflection $s$ across the hyperplane perpendicular to the fifth lattice basis vector $v_{4}$ shown in Table 4.2. The reflection $s$ is the same reflection as in equation (2.1), and it corresponds to a reflection across a horizontal line in Figure 4.2. Let $H$ be the subgroup of $G$ generated by $S_{5}$ and $s$ (the symmetry group of the kissing configuration $Q_{5})$, let $H_{0}=H$, and let $H_{1}=(-1)H$ be the coset obtained through scalar multiplication by $-1$. Then a case analysis shows that the symmetry group of the $Q_{5}$ uniform packing is
@@
For the $R_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{3}$ root lattice and two differently scaled one-dimensional lattices (again see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{4}\times C_{2}^{3}$, where the symmetric group $S_{4}$ permutes the first four coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by $-1$ and the reflections $r$ and $s$ across the hyperplanes perpendicular to the fourth and fifth lattice basis vectors $v_{3}$ and $v_{4}$ shown in Table 4.2. The reflections $r$ and $s$ correspond to reflections across vertical and horizontal lines in Figure 4.2, respectively. Let $H$ be the subgroup of $G$ generated by $S_{4}$ and $s$ (the symmetry group of the kissing configuration $R_{5})$, and let $H_{0}=H$, $H_{1}=(-r)H$, $H_{2}=rH$, and $H_{3}=(-1)H$ be its cosets, where as above $(-1)$ denotes scalar multiplication by $-1$ and $-r$ denotes the composition of $(-1)$ and $r$. Then a case analysis shows that the symmetry group of the $R_{5}$ uniform packing is
@@ 4.4. Higher dimensions
Matching the density of the Leech lattice would amount to achieving $\delta=2^{13}=8192$, which turns out to be impossible because each of the four colors gives a sphere packing with radius 2/$\sqrt{2}/2$, and achieving $\delta=8192$ would require impossibly dense sphere packings. Specifically, the Blichfeldt bound implies that each color has at most $788.5785$ points per unit volume (see, for example, Theorem 6.1 in [23]), and four times that bound is still far below $8192$.

### v3

@@ 1. Introduction
Figure 1.1. The centers of the spheres in the face-centered cubic kissing arrangement on the left, and the centers of the spheres in the hexagonal close packing kissing arrangement on the right. To produce the arrangement on the right from the one on the left, the top layer is deleted and the bottom layer is reflected across the central layer (the hexagon).
@@ 2. Five-dimensional kissing configurations
The kissing number in five dimensions appears to be $40$, although the best upper bound that has been proved is $44$ (from [15]). The first construction achieving $40$ is implicit in Korkine and Zolotareff’s 1873 paper [9], where they constructed the $D_{5}$ root lattice. Its root system achieves a kissing number of $40$ as the permutations of the points $(\pm 1,\pm 1,0,0,0)$; these points form a kissing configuration because they each have squared norm $2$ and the inner product between distinct points is always at most $1$.
In 1967, Leech [11] constructed a different kissing configuration of the same size, not isometric to the $D_{5}$ root system. He split $D_{5}$ into four-dimensional layers based on the fifth coordinate: there are $8$ points with fifth coordinate $1$, $24$ with 0, and $8$ with 1-1. The top layer consists of the points
$(\overbrace{\pm 1,0,0,0}^{\text{permute}},1),$
@@
$(\overbrace{\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2},\pm\tfrac{1}{2}}^{\text{odd \# of $-$ signs}},1)$
yields a different kissing configuration $L_{5}$, which breaks the reflection symmetry across the central hyperplane with fifth coordinate 0. One could also use an even number of minus signs, or replace both the top and bottom layers, but these modifications are all isometric to $D_{5}$ or $L_{5}$. For comparison, such a construction cannot be carried out in four dimensions, since the optimal kissing configuration is unique in four dimensions [6].
@@
$(\overbrace{\pm 1,\pm 1,0,0}^{\text{permute}},0)$
@@
[…]
@@
cyclically permutes these cross polytopes in four dimensions, so they all play the same role. Up to isometry, there are only two kissing configurations that can be obtained in this way: $D_{5}$, which has reflection symmetry across the hyperplane, and $L_{5}$, which does not. One can see that they are distinct by counting antipodal points: $D_{5}$ is closed under multiplication by 1-1, while $L_{5}$ is not. Instead, only $24$ of the points in $L_{5}$ come in antipodal pairs, namely those in $D_{4}$.
In 2023, Szöllősi [20] found a new kissing configuration $Q_{5}$ via a computer search. It can be described similarly to $L_{5}$, but using different cross sections: the $D_{5}$ root system contains $20$ vectors with a coordinate sum of 0 and $10$ each with coordinate sums of $\pm 2$. Here the central cross section is the $A_{4}$ root system. This cross section is suboptimal as a four-dimensional kissing configuration, but it nevertheless extends to a seemingly optimal configuration in five dimensions. The layer with coordinate sum $2$ consists of the antipodes of the points with coordinate sum 2-2, but it is not the reflection of that layer across the hyperplane with coordinate sum 0. Reflection across this hyperplane consists of the map
@@
where $w=(1,1,1,1,1)/\sqrt{5}$ and $\langle\cdot,\cdot\rangle$ denotes the usual inner product. Equivalently, when applied to a vector with coordinate sum $s$, the reflection subtracts $2s/5$ from each coordinate. The $Q_{5}$ kissing configuration replaces the layer that has coordinate sum $2$ with the reflection of the layer with coordinate sum 2-2. Only $20$ points in $Q_{5}$ come in antipodal pairs, namely those in $A_{4}$, and so it cannot be isometric to $D_{5}$ or $L_{5}$.
Our new configuration, which we call $R_{5}$, can be obtained as follows by modifying $L_{5}$. The $L_{5}$ configuration behaves similarly to $D_{5}$ with respect to coordinate sums: there are $20$ points with coordinate sum 0 and $10$ each with $\pm 2$. The central cross section is no longer the $A_{4}$ root system, but rather a modification that replaces four adjacent points with a different regular tetrahedron. Again we can replace one of the adjacent layers with the other’s reflection, to obtain $R_{5}$ from $L_{5}$. Only $12$ of the points in $R_{5}$ come in antipodal pairs, and so it cannot be isometric to any of the three other configurations. This yields the following theorem:
@@
Configuration 1-1 4/5-4/5 3/4-3/4 1/2-1/2 3/10-3/10 1/4-1/4 0 $1/5$ $1/2$
$D_{5}$ 2020 0 0 $240$ 0 0 $280$ 0 $240$
$L_{5}$ $12$ 0 $32$ $192$ 0 $32$ $272$ 0 $240$
$Q_{5}$ $10$ $30$ 0 $180$ $60$ 0 $250$ $10$ $240$
$R_{5}$ 66 $30$ 2020 $144$ $60$ $28$ $242$ $10$ $240$
Table $2.$ The four five-dimensional kissing configurations with $40$ points. The subdivisions in the list are $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$, respectively, and the points with fractional coordinates are the only ones that differ from those in $D_{5}$.
@@
Table 2.1 lists the inner products that occur in these configurations and their multiplicities, while Table 2.2 lists coordinates for the points themselves. The symmetry groups of $D_{5}$, $L_{5}$, $Q_{5}$, and $R_{5}$ have orders $3840$, $384$, $240$, and $48$, respectively. Specifically, the symmetries of $D_{5}$ are the signed permutations of the coordinates (equivalently, the $1920$ elements of the $D_{5}$ Weyl group times two Dynkin diagram automorphisms), and those of $L_{5}$ are the signed permutations that fix the fifth coordinate. The symmetries of $Q_{5}$ are generated by the permutations of the coordinates and reflection across the hyperplane, and the symmetries of $R_{5}$ are generated by the permutations of the coordinates that fix the fifth coordinate and reflection across the hyperplane.
@@ 3. Six-dimensional kissing configurations
$(\pm\sqrt{2/5},\dots,\pm\sqrt{2/5})$
@@
If we start with $Q_{5}$ or $R_{5}$, it initially seems that the same approach might produce an extension to six dimensions, but it cannot be fully carried out. These configurations again have $32$ deep holes: those on one side of the hyperplane with coordinate sum 0 are the same as in $D_{5}$ or $L_{5}$, respectively, while those on the other side are the reflections across this hyperplane. However, continuing the construction requires a set of $16$ holes with maximal inner product $2/5$ between them, and one can check by depth-first search that no such set exists. We have not ruled out the possibility that some other construction might give a six-dimensional kissing configuration with $Q_{5}$ or $R_{5}$ as a cross section, but it seems unlikely:
@@ 4.1. Packings that fiber over $D_{3}$
To form a five-dimensional packing, we will use a four-colored point configuration in $\mathbb{R}^{2}$, by which we mean an infinite discrete point set in $\mathbb{R}^{2}$ in which each point is labeled with the color 0, 11, 22, or $3$. We call two colors adjacent if they differ by $1$ modulo $4$, and opposite if they differ by $2$. If the point $v$ is colored $i$, then we center spheres of radius $\sqrt{2}/2$ at the points in $\{v\}\times(D_{3}+t_{i})$. In order for the spheres not to overlap, the four-colored point configuration must satisfy the following constraints:
@@
A valid coloring is one that satisfies these constraints. If a valid four-colored point configuration has $\delta$ points per unit area in $\mathbb{R}^{2}$, then the corresponding sphere packing in $\mathbb{R}^{5}$ has packing density given by
$\frac{\pi^{5/2}}{(5/2)!}\left(\frac{\sqrt{2}}{2}\right)^{5}\cdot\delta\cdot\frac{1}{2}=\frac{\pi^{2}\sqrt{2}}{30}\,\delta.$
Here, we are multiplying the volume of a sphere of radius $\sqrt{2}/2$ in $\mathbb{R}^{5}$ with the point densities of the four-colored point configuration in $\mathbb{R}^{2}$ and the $D_{3}$ root lattice in $\mathbb{R}^{3}$. In particular, matching the packing density of the $D_{5}$ root lattice amounts to achieving $\delta=1$. Presumably it is impossible to achieve $\delta>1$:
@@
To achieve $\delta=1$, we can tile the plane with two types of triangles: isosceles triangles with side lengths $\sqrt{5}/2,\sqrt{5}/2,1$ or $1,1,\sqrt{2}$, which we call $\Delta_{1}$ and $\Delta_{2}$. Both of these triangles have area $1/2$, and it follows that in every edge-to-edge tiling, there is one vertex per unit area. [^2] The angles in a triangle add up to $\pi$, which means they account for half of the total angle surrounding a single vertex. Therefore there must be half a vertex per triangle on average in any edge-to-edge tiling. Our five-dimensional sphere packings are obtained from valid four-colorings of the vertices of these tilings. Specifically, vertices at distance $\sqrt{2}$ have the same color, those at distance $\sqrt{5}/2$ have adjacent colors, and those at distance $1$ have opposite colors.
[…]
@@
As mentioned in the introduction, the existence of more general packings such as those in Figure 4.2 disproves conjectures of Conway and Sloane [5] and Kuperberg [10]. We propose the following conjecture as a salvage for Conjecture 15 in [10].
@@ 4.2. Uniform packings
The uniform packing from Figure 4.2 with kissing configuration $Q_{5}$ is $2$-periodic, and the one with kissing configuration $R_{5}$ is $4$-periodic, with lattice bases and translation vectors shown in Table 4.2. Andreanov and Kallus [1] showed that no $2$-periodic sphere packing in $\mathbb{R}^{5}$ can be denser than the $D_{5}$ root lattice, and they obtained a complete list of the densest such packings. In the process they discovered the $Q_{5}$ uniform sphere packing, which also occurs in the acknowledgements of [20]. No such optimality theorem or classification is known beyond $2$-periodic packings. As for the remaining uniform packings, the $D_{5}$ root lattice is of course $1$-periodic, and the three $L_{5}$ uniform packings are $2$-periodic, 33-periodic, and $4$-periodic (see [5]).
@@ 4.3. Symmetries of uniform packings
For the $Q_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{4}$ root lattice and a one-dimensional lattice (see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{5}\times C_{2}^{2}$, where the symmetric group $S_{5}$ permutes the five coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by 1-1 and the reflection $s$ across the hyperplane perpendicular to the fifth lattice basis vector $v_{4}$ shown in Table 4.2. The reflection $s$ is the same reflection as in equation (2.1), and it corresponds to a reflection across a horizontal line in Figure 4.2. Let $H$ be the subgroup of $G$ generated by $S_{5}$ and $s$ (the symmetry group of the kissing configuration $Q_{5})$, let $H_{0}=H$, and let $H_{1}=(-1)H$ be the coset obtained through scalar multiplication by 1-1. Then a case analysis shows that the symmetry group of the $Q_{5}$ uniform packing is
@@
For the $R_{5}$ uniform packing, the underlying lattice $\Lambda$ is the orthogonal direct sum of the $A_{3}$ root lattice and two differently scaled one-dimensional lattices (again see Table 4.2 for a lattice basis). The symmetry group $G$ of $\Lambda$ is isomorphic to $S_{4}\times C_{2}^{3}$, where the symmetric group $S_{4}$ permutes the first four coordinates and the cyclic factors $C_{2}$ are generated by scalar multiplication by 1-1 and the reflections $r$ and $s$ across the hyperplanes perpendicular to the fourth and fifth lattice basis vectors $v_{3}$ and $v_{4}$ shown in Table 4.2. The reflections $r$ and $s$ correspond to reflections across vertical and horizontal lines in Figure 4.2, respectively. Let $H$ be the subgroup of $G$ generated by $S_{4}$ and $s$ (the symmetry group of the kissing configuration $R_{5})$, and let $H_{0}=H$, $H_{1}=(-1)H$, $H_{2}=rH$, and $H_{3}=(-r)H$ be its cosets, where as above (1)(-1) denotes scalar multiplication by 1-1 and (r)(-r) denotes the composition of (1)(-1) and $r$. Then a case analysis shows that the symmetry group of the $R_{5}$ uniform packing is
@@ 4.4. Higher dimensions
Matching the density of the Leech lattice would amount to achieving $\delta=2^{13}=8192$, which turns out to be impossible because each of the four colors gives a sphere packing with radius $\sqrt{2}/2$, and achieving $\delta=8192$ would require impossibly dense sphere packings. Specifically, the Blichfeldt bound implies that each color has at most $788.5785$ points per unit volume (see, for example, Theorem 6.1 in [23]), and four times that bound is still far below $8192$.